# Efficient Diffusion Models: A Survey

### v1

@@ Abstract
Diffusion models have emerged as powerful generative models capable of producing high-quality contents such as images, videos, and audio, demonstrating their potential to revolutionize digital content creation. However, these capabilities come at the cost of their significant computational resources and lengthy generation time, underscoring the critical need to develop efficient techniques for practical deployment. In this survey, we provide a systematic and comprehensive review of research on efficient diffusion models. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient diffusion model topics from algorithm-level, system-level, and framework perspective, respectively. We have also created a GitHub repository where we organize the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-Diffusion-Model-Survey. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of efficient diffusion model research and inspire them to contribute to this important and exciting field.
@@ 1 Introduction
Diffusion models kickstart a new era in the field of artificial intelligence generative content (AIGC), garnering unprecedented attention (Yang et al., 2023b; Croitoru et al., 2023b). Especially in the context of image synthesis tasks, diffusion models have demonstrated impressive and diverse generative capabilities. The powerful cross-modal capabilities of diffusion models have also further fueled the vigorous development of downstream tasks (Chen et al., 2023b). Despite the increasing maturity of diffusion model variants after numerous iterations (Zhang et al., 2023d; Xu et al., 2023), generating high-resolution complex natural scenes remains both time-consuming and computationally intensive, whether the initial pixel-level approach (Ho et al., 2020) or the latent space variant (Rombach et al., 2022). Therefore, in order to optimize user-level deployment of diffusion models, researchers have never ceased their pursuit of efficient diffusion models.
@@
While there are numerous comprehensive surveys on diffusion models (Yang et al., 2023b; Chen et al., 2024; Croitoru et al., 2023a; Cao et al., 2024) and those focused on specific fields and tasks (Ulhaq et al., 2022; Lin et al., 2024c; Kazerouni et al., 2023; Lin et al., 2024b; Peng et al., 2024b; Daras et al., 2024), discussions on the efficiency of diffusion models are notably scarce. The only existing survey addressing efficiency (Ma et al., 2024c) serves as an initial exploration in this area. In our work, we provide a more comprehensive and detailed taxonomy of efficient techniques, covering a broader and more recent collection of research papers.
The overarching goal of this survey is to provide a holistic view of the technological advances in efficient diffusion models from algorithm-level, system-level, application, and framework perspectives, as illustrated in Figure 1. These four categories cover distinct yet interconnected research topics, collectively providing a systematic and comprehensive review of efficient diffusion models research. Specifically,
@@
  • Applications: Applications of diffusion models span various domains, where efficiency directly impacts their practical usability. Tailored techniques are required to optimize the performance of diffusion models for these specific tasks without sacrificing quality. In §2.2, we survey efficient techniques that focus on image generation, video generation, text generation, audio generation, and 3D generation.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em, font=, where level=2text width=18em, font=, where level=3text width=16em, font=, where level=4text width=18em, font=, where level=5text width=18em, font=, [ Efficient Diffusion, ver [ Application (§2.2) , fill=green!10 [ Image Generation (§2.2.1), fill=green!10 ] [ Video Generation (§2.2.2), fill=green!10 ] [ Text Generation (§2.2.3), fill=green!10 ] [ Audio Generation (§2.2.4), fill=green!10 ] [ 3D Generation (§2.2.5), fill=green!10 ] ] [ Algorithm (§3), fill=blue!10 [ Efficient Training (§3.1), fill=blue!10 [ Noise Schedule, fill=blue!10 ] [ Score Matching, fill=blue!10 ] [ Data-Dependent Adaptive Priors , fill=blue!10 ] [ Rectified Flow, fill=blue!10 ] ] [ Efficient Fine-tuning (§3.2), fill=blue!10 [ LoRA, fill=blue!10 ] [ Adapter, fill=blue!10 ] [ ControlNet, fill=blue!10 ] ] [ Efficient Sampling (§3.3), fill=blue!10 [ Efficient Solver, fill=blue!10 [ SDE Solver, fill=blue!10 ] [ ODE Solver, fill=blue!10 ] ] [ Sampling Scheduling, fill=blue!10 [ Parallel Sampling, fill=blue!10 ] [ Timestep Schedule, fill=blue!10 ] ] [ Partial Sampling, fill=blue!10 [ Early Exit, fill=blue!10 ] [ Retrieval-Based Diffusion, fill=blue!10 ] ] ] [ Compression (§3.4), fill=blue!10 [ Quantization, fill=blue!10 [ Post-Training Quantization, fill=blue!10 ] [ Quantization-Aware Training, fill=blue!10 ] ] [ Pruning, fill=blue!10 ] [ Knowledge Distillation, fill=blue!10 [ Vector Field Distillation, fill=blue!10 ] [ Generator Distillation, fill=blue!10 ] ] ] ] [ System (§4), fill=orange!10 [ Hardware-Software Co-Design (§4.1) , fill=orange!10 ] [ Parallel Computing (§4.2) , fill=orange!10 ] [ Caching Technique (§4.3) , fill=orange!10 ] ] [ Frameworks (§5), fill=lime!10 [ Diffusers, DALL-E, OneDiff, LiteGen, InvokeAI, Comfy UI-Docker, Grate, DifFUSER,
Versatile Diffusion, UniDiffuser, text width=43em, fill=lime!10 ] ] ]
@@ 2 Background and Applications
2.1 Background of Diffusion Models
@@
2.2 Applications of Diffusion Models
@@ 2.2.2 Video Generation
Following the rapid escalation in image generation, video generation similarly garnered widespread attention (Ho et al., 2022; Singer et al., 2022; Rombach et al., 2022; Xing et al., 2023). The heavy model size and the substantial computational costs have further intensified the focus on developing more efficient methods for video generation (Zhang et al., 2023a; Liu et al., 2023a; Xing et al., 2024; Wang et al., 2023; Lee et al., 2024b). For example, Zhang et al. (2023a) introduced AdaDiff, a lightweight framework designed to optimize a specialized policy gradient method tailored to individual text prompts. This approach facilitates the design of reward functions and enables an effective trade-off between inference time and generation quality. Specifically to the training process, Liu et al. (2023a) proposed an efficient training framework ED-T2V to freeze the pretraining model (Rombach et al., 2022) training additional temporal modules. Similarly, Xing et al. (2024) suggested using spatial and temporal adapters. In their approach, the original T2I model remains frozen during training, and only the newly added adapter modules are updated. Unlike the works above, Wang et al. (2023) presented VideoLCM, incorporating consistency distillation in the latent space. VideoLCM efficiently distills knowledge from a pretraining model, maintaining fidelity and temporal coherence while improving inference speed.  Lee et al. (2024b) introduces a grid diffusion model by representing a video as a grid of images. It employs key grid image generation and autoregressive grid interpolation to maintain temporal consistency.
@@ 2.2.4 Audio Generation
In the field of audio generation, the application of diffusion models presents several unique challenges. First, audio data exhibits strong temporal continuity, especially in high-resolution audio generation tasks, where the model must accurately reconstruct both time-domain and frequency-domain information. Compared to images or text, even subtle distortions or noise in audio are easily perceptible by humans, directly affecting the listening experience, particularly in speech and music generation tasks. Ensuring high fidelity and maintaining the consistency of details in the generated audio is therefore crucial. Moreover, many audio generation tasks require low-latency feedback, especially in applications like speech synthesis and real-time dialogue, which makes acceleration of diffusion models essential. The multi-dimensional nature of audio data, such as time-domain, frequency-domain, stereo, and spatial audio, further complicates the generation process, requiring the model to handle these dimensions while maintaining consistency during the accelerated generation. To address these challenges, researchers have proposed various methods to accelerate diffusion models in audio generation. Some works focus on reducing the number of diffusion steps to speed up the generation process, such as Chen et al. (2020) in WaveGrad and Kong et al. (2020) in DiffWave, which optimize the network structure to reduce generation time while maintaining high audio quality. Further optimization comes from the FastDPM framework (Kong & Ping, 2021), which generalizes discrete diffusion steps to continuous ones, using a bijective mapping between noise levels to accelerate sampling without compromising quality. FastDPM’s flexibility allows it to adapt to different domains, and in the case of audio synthesis, where stochasticity plays a crucial role, it demonstrates superior performance in high-stochasticity tasks like speech generation. Through these approaches, diffusion models not only accelerate the generation
@@
process but also reduce computational costs while ensuring that audio quality remains high, meeting the demands of real-time audio generation applications.
2.2.5 3D Generation
As a technique closely aligned with real-world representation, 3D generation holds substantial promise across various sectors, including medical imaging, motion capture, asset production, and scene reconstruction, etc. However, when compared to 2D image generation, distinctive high-resolution elements such as volumetric data or point clouds present unique challenges, significantly escalating computational demands. Several efficient methodologies (Bieder et al., 2023; Zhou et al., ; Tang et al., 2023; Park et al., 2023) have been proposed, particularly concentrating on enhancing the sampling process and optimizing the architectural framework, which further handles the computational complexity inherent. One of the most prevalent approaches involves designing more efficient sampling schedules (Bieder et al., 2023; Li et al., 2024c; Yu et al., 2024; Zhou et al., ). By utilizing larger sampling step sizes, modifying the sampling strategy between 2D and 3D, or incorporating multi-view parallelism, these techniques address the key bottlenecks in the sampling process, thereby improving sampling efficiency. Moreover, the incorporation of novel architectures, such as state-space models and lightweight feature extractors (Mo, 2024; Tang et al., 2023), alleviates the computational burden of processing 3D data, significantly enhancing model efficiency.

### v2

@@ Abstract
Diffusion models have emerged as powerful generative models capable of producing high-quality contents such as images, videos, audio, 3D, and text demonstrating their potential to revolutionize digital content creation. However, these capabilities come at the cost of their significant computational resources and lengthy generation time, underscoring the critical need to develop efficient techniques for practical deployment. In this survey, we provide a systematic and comprehensive review of research on efficient diffusion models. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient diffusion model topics from algorithm-level, system-level, and framework perspective, respectively. We have also created a GitHub repository where we organize the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-Diffusion-Model-Survey. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of efficient diffusion model research and inspire them to contribute to this important and exciting field.
@@ 1 Introduction
Diffusion models kickstart a new era in the field of artificial intelligence generative content (AIGC), garnering unprecedented attention (Yang et al., 2023b; Croitoru et al., 2023). Especially in the context of image synthesis tasks, diffusion models have demonstrated impressive and diverse generative capabilities. The powerful cross-modal capabilities of diffusion models have also further fueled the vigorous development of downstream tasks (Chen et al., 2023b). Despite the increasing maturity of diffusion model variants after numerous iterations (Zhang et al., 2023d; Xu et al., 2023a), generating high-resolution complex natural scenes remains both time-consuming and computationally intensive, whether the initial pixel-level approach (Ho et al., 2020) or the latent space variant (Rombach et al., 2022). Therefore, in order to optimize user-level deployment of diffusion models, researchers have never ceased their pursuit of efficient diffusion models.
@@
While there are numerous comprehensive surveys on diffusion models (Yang et al., 2023b; Chen et al., 2024; Croitoru et al., 2023; Cao et al., 2024) and those focused on specific fields and tasks (Ulhaq et al., 2022; Lin et al., 2024c; Kazerouni et al., 2023; Lin et al., 2024b; Peng et al., 2024b; Daras et al., 2024), discussions on the efficiency of diffusion models are notably scarce. The only existing survey addressing efficiency (Ma et al., 2024c) serves as an initial exploration in this area. In our work, we provide a more comprehensive and detailed taxonomy of efficient techniques, covering a broader and more recent collection of research papers.
The overarching goal of this survey is to provide a holistic view of the technological advances in efficient diffusion models from algorithm-level, system-level, and framework perspectives, as illustrated in Figure 1. These four categories cover distinct yet interconnected research topics, collectively providing a systematic and comprehensive review of efficient diffusion models research. Specifically,
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em, font=, where level=2text width=18em, font=, where level=3text width=16em, font=, where level=4text width=16em, font=, where level=5text width=18em, font=, [ Efficient Diffusion, ver [ Algorithm (§3), fill=blue!10 [ Efficient Training (§3.1), fill=blue!10 [ Latent Diffusion, fill=blue!10 ] [ Loss Formulation, fill=blue!10 [ Score Matching, fill=blue!10 ] [ Rectified Flow, fill=blue!10 ] ] [ Training Tricks , fill=blue!10 [ Data-Dependent Adaptive Priors , fill=blue!10 ] [ Extra Loss Supervision , fill=blue!10 ] [ Noise Schedule , fill=blue!10 ] ] ] [ Efficient Fine-tuning (§3.2), fill=blue!10 [ LoRA, fill=blue!10 ] [ Adapter, fill=blue!10 ] [ ControlNet, fill=blue!10 ] ] [ Efficient Sampling (§3.3), fill=blue!10 [ Efficient Solver, fill=blue!10 [ SDE Solver, fill=blue!10 ] [ ODE Solver, fill=blue!10 ] ] [ Sampling Scheduling, fill=blue!10 [ Parallel Sampling, fill=blue!10 ] [ Timestep Schedule, fill=blue!10 ] ] [ Truncated Sampling, fill=blue!10 [ Early Exit, fill=blue!10 ] [ Retrieval-Based Diffusion, fill=blue!10 ] ] [ Knowledge Distillation, fill=blue!10 [ Vector Field Distillation, fill=blue!10 ] [ Generator Distillation, fill=blue!10 ] ] ] [ Compression (§3.4), fill=blue!10 [ Quantization, fill=blue!10 [ Post-Training Quantization, fill=blue!10 ] [ Quantization-Aware Training, fill=blue!10 ] ] [ Pruning, fill=blue!10 ] ] ] [ System (§4), fill=orange!10 [ Hardware-Software Co-Design (§4.1) , fill=orange!10 ] [ Parallel Computing (§4.2) , fill=orange!10 ] [ Caching Technique (§4.3) , fill=orange!10 ] ] [ Frameworks(§5), fill=lime!10 [ Flash Attention, xFormers, DeepSpeed, OneFlow, Stable-Fast, Onediff, DeepCache, TGATE, xDiT, text width=49em, fill=lime!10 ] ] ]
@@ 2 Background and Applications
2.1 Basic formulas for Diffusion models
Diffusion models have emerged as powerful generative models capable of producing high-quality samples across various domains. This section explores the theoretical foundations and recent advancements in the diffusion model framework. We begin by examining the fundamental formulation of Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), which leverage forward and backward stochastic processes to gradually transform data into noise and then reverse this process for generation. We then explore Score Matching (Hyvärinen & Dayan, 2005) as an alternative formulation that directly optimizes gradient fields of probability densities. Continuing our discussion, we analyze how Stochastic Differential Equations (SDEs) and Ordinary Differential Equations (ODEs) (Song et al., 2020c) provide continuous perspectives on diffusion. Finally, we introduce Flow Matching (Lipman et al., 2022) as a recent paradigm that offers improved efficiency by directly learning vector fields that transform distributions, utilizing ODEs to provide a deterministic process.
2.1.1 DDPM
@@
2.1.2 Score Matching
While DDPM directly optimizes the noise prediction in Eq.(6), score matching objectives can directly be estimated on a dataset and optimized with stochastic gradient descent. The loss function for score matching takes a different approach, formulated as follows:
2.1.3 Solvers
Given that the cost of sampling escalates proportionally with the number of discretized time steps, many researchers have concentrated on devising discretization schemes that reduce the number of time steps. A key insight emerges from reexamining the discrete forward process in the original DDPM formulation Eq.(1), as we reduce the step size between consecutive steps, the process naturally approaches a continuous transformation. Consequently, adopting learning-free methods using SDE or ODE solvers (Song et al., 2020c) has been proposed.
SDE Solver. Song et al. (2020c) firstly presents a stochastic differential equation (SDE) that smoothly transforms a complex data distribution to a known prior distribution by slowly injecting noise and a corresponding reverse-time SDE that transforms the prior distribution back into the data distribution by slowly removing the noise.The discrete noise addition steps in Eq.(1) are reformulated into a continuous process:
$\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\,\mathrm{d}t+g(t)\,\mathrm{d}\bar{\mathbf{w}}$ (10)
$\mathrm{d}\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\right]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}$ (11)
ODE Solver. Unlike SDE solvers, the trajectories generated by ordinary differential equation (ODE) solvers are deterministic (Song et al., 2020c), remaining unaffected by stochastic variations. Consequently, these deterministic ODE solvers tend to achieve convergence more rapidly compared to their stochastic counterparts, although this often comes at the expense of a marginal reduction in sample quality. The corresponding deterministic process Eq.(12) can be derived from the reverse-time SDE Eq.(11) by removing the stochastic term $g(t)\mathrm{d}\bar{\mathbf{w}}$, resulting in a deterministic process that shares the same marginal probability densities as the reverse-SDE:
$\mathrm{d}\mathbf{x}=\left[{\mathbf{f}(\mathbf{x},t)}-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})\right]\mathrm{d}t$ (12)
$\mathrm{d}\mathbf{x}=\mathbf{f}(\mathbf{x},t)\mathrm{d}t$ (13)
2.1.4 Flow Matching
Flow Matching (FM) (Lipman et al., 2022) is a new paradigm for generative modeling based on Continuous Normalizing Flows (CNFs). This approach allows us to train CNFs at an unprecedented scale and offers greater efficiency compared to traditional diffusion models.
To better understand the process of flow matching, we need to dive into the preliminaries. First of all, we make $\mathbb{R}^{d}$ denote the data space with data points $x=(x^{1},\ldots,x^{d})\in\mathbb{R}^{d}$. Then, we need to understand several key concepts: Probability density path $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$, which is a time-dependent probability density function, and Time-dependent vector field $v_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, which describes how data points change over time. A vector field $v_{t}$ can be used to construct a time-dependent diffeomorphic map, defined by an ordinary differential equation (ODE):
$\frac{\mathrm{d}}{\mathrm{d}t}\phi_{t}(x)=v_{t}(\phi_{t}(x))$ (14)
$\phi_{0}(x)=x$ (15)
A Continuous Normalizing Flow (CNF) is a generative model that parametrizes the time-dependent vector field $v_{t}(x;\theta)$ using a neural network (Chen et al., 2018), where $\theta$ represents the learnable parameters. This vector field defines a flow $\phi_{t}$ that transforms the probability distribution from a simple prior density $p_{0}$ to a more complex target density $p_{1}$ through the push-forward equation. The core of CNF is using this flow to reshape probability distributions:
$p_{t}=[\phi_{t}]_{*}p_{0}$ (16)
The push-forward operator * is defined as:
$[\phi_{t}]_{*}p_{0}(x)=p_{0}(\phi_{t}^{-1}(x))\det\left[\frac{\partial\phi_{t}^{-1}}{\partial x}(x)\right]$ (17)
The FM loss is formally defined as:
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x\sim p_{t}(x)}\left\|v_{t}(x)-u_{t}(x)\right\|^{2},$ (18)
where: $u_{t}(x)$ generates a predefined probability density path $p_{t}(x)$, $p_{0}(x)=\mathcal{N}(x|0,I)$ is a simple prior distribution (e.g., Gaussian noise), and $p_{1}(x)\approx q(x)$ approximates the data distribution.
While FM is conceptually straightforward, directly optimizing this objective is intractable due to the lack of closed-form expressions for $p_{t}$ and $u_{t}$. To address this, FM constructs conditional probability paths $p_{t}(x|x_{1})$ and vector fields $u_{t}(x|x_{1})$ per data sample $x_{1}\sim q(x_{1})$, where $p_{0}(x|x_{1})=p(x)$ at time $t=0$, and $p_{1}(x|x_{1})=\mathcal{N}(x|x_{1},\sigma^{2}I)$ at $t=1$ (a normal distribution with $x_{1}$ mean and a sufficiently small standard deviation $\sigma>0$). These conditional paths are then aggregated into global counterparts through marginalization:
$p_{t}(x)=\int p_{t}(x|x_{1})q(x_{1})\mathrm{d}x_{1}$ (19)
This approach avoids explicit modeling of the intractable marginal distributions and enables scalable training, while ensuring that at $t=1$, the marginal distribution $p_{1}(x)$ approximates the data distribution $q(x)$.
The marginal vector field can then be defined as:
$u_{t}(x)=\int u_{t}(x|x_{1})\frac{p_{t}(x|x_{1})q(x_{1})}{p_{t}(x)}\mathrm{d}x_{1}$ (20)
A key theorem in the paper shows that this constructed marginal vector field $u_{t}$ in Eq.(20) indeed generates the marginal probability path $p_{t}$ in Eq.(19).
However, computing the FM objective is still infeasible because it involves complex integrals. To address this issue, the paper proposes the Conditional Flow Matching (CFM) objective:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,q(x_{1}),p_{t}(x|x_{1})}\|v_{t}(x)-u_{t}(x|x_{1})\|^{2},$ (21)
The second key theorem in the paper proves that the FM and CFM objectives have identical gradients with respect to $\theta$, so optimizing CFM is equivalent to optimizing FM, but computationally simpler. Compared to score matching, flow matching is faster because score matching typically requires computing Hessian matrices in Eq(9), while CFM cleverly avoids such complex calculations by only computing the L2 distance between vector fields. In conclusion, it’s worth noting that Flow Matching can be theoretically connected with other diffusion model formulations. As summarized by Gao et al. (2024), flow matching, score-based models, and traditional diffusion models can be understood within a unified mathematical framework. These approaches essentially represent different parameterizations of the same continuous-time process. While traditional diffusion models like DDPM (Ho et al., 2020) learn to estimate noise or score functions, and score-based models directly optimize gradient fields (Hyvärinen & Dayan, 2005), flow matching learns the vector field itself, directly capturing the transformation map between distributions.
2.2 Applications
Building on the foundational principles of diffusion models outlined in  §2.1, this section surveys their practical deployment across diverse generative tasks, with a specific emphasis on efficiency-driven innovations. As computational demands and real-world applicability become increasingly critical, researchers have adapted diffusion models to optimize resource usage while preserving or enhancing output quality. Here, we explore key application domains—image, video, text, audio, and 3D generation—highlighting techniques that reduce inference time, memory footprint, and training complexity. These advancements underscore the adaptability of diffusion models and their growing impact in addressing the challenges of scalable, high-fidelity content generation.
@@ 2.2.2 Video Generation
Following the rapid escalation in image generation, video generation similarly garnered widespread attention (Melnik et al., 2024; Ho et al., 2022b; Xing et al., 2023). The heavy model size and the substantial computational costs have further intensified the focus on developing more efficient methods for video generation  (Zhang et al., 2023b; Liu et al., 2023b; Xing et al., 2024; Wang et al., 2023; Lee et al., 2024b). For example, Zhang et al. (2023b) introduced AdaDiff, a lightweight framework designed to optimize a specialized policy gradient method tailored to individual text prompts. This approach facilitates the design of reward functions and enables an effective trade-off between inference time and generation quality. Specifically to the training process, Liu et al. (2023b) proposed an efficient training framework ED-T2V to freeze the pretraining model (Rombach et al., 2022) training additional temporal modules. Similarly, Xing et al. (2024) suggested using spatial and temporal adapters. In their approach, the original T2I model remains frozen during training, and only the newly added adapter modules are updated. Unlike the works above, Wang et al. (2023) presented VideoLCM, incorporating consistency distillation in the latent space. VideoLCM efficiently distills knowledge from a pretraining model, maintaining fidelity and temporal coherence while improving inference speed.  Lee et al. (2024b) introduces a grid diffusion model by representing a video as a grid of images. It employs key grid image generation and autoregressive grid interpolation to maintain temporal consistency. Moreover,  Ceylan et al. (2023) leverages self-attention feature injection and guided latent updates, efficiently repurposing image models for video editing, enabling high-quality, consistent edits at minimal computational overhead.  Yin et al. (2023) proposes NUMA-XL, using a hierarchical coarse-to-fine approach to enable parallelizable, exponential scaling of video length. It achieves a great reduction in inference time and is trained directly on long sequences, ensuring long-term consistency.
@@ 2.2.4 Audio Generation
In the field of audio generation, the application of diffusion models presents several unique challenges. First, audio data exhibits strong temporal continuity, especially in high-resolution audio generation tasks, where the model must accurately reconstruct both time-domain and frequency-domain information. Compared to images or text, even subtle distortions or noise in audio are easily perceptible by humans, directly affecting the listening experience, particularly in speech and music generation tasks. Ensuring high fidelity and maintaining the consistency of details in the generated audio is therefore crucial. Moreover, many audio generation tasks require low-latency feedback, especially in applications like speech synthesis and real-time dialogue, which makes acceleration of diffusion models essential. The multi-dimensional nature of audio data, such as time-domain, frequency-domain, stereo, and spatial audio, further complicates the generation process, requiring the model to handle these dimensions while maintaining consistency during the accelerated generation. To address these challenges, researchers have proposed various methods to accelerate diffusion models in audio generation. Some works focus on reducing the number of diffusion steps to speed up the generation process, such as Chen et al. (2020) in WaveGrad and Kong et al. (2020) in DiffWave, which optimize the network structure to reduce generation time while maintaining high audio quality. Further optimization comes from the FastDPM framework (Kong & Ping, 2021), which generalizes discrete diffusion steps to continuous ones, using a bijective mapping between noise levels to accelerate sampling without compromising quality. FastDPM’s flexibility allows it to adapt to different domains, and in the case of audio synthesis, where stochasticity plays a crucial role, it demonstrates superior performance in high-stochasticity tasks like speech generation. Through these approaches, diffusion models not only accelerate the generation process but also reduce computational costs while ensuring that audio quality remains high, meeting the demands of real-time audio generation applications.
2.2.5 3D Generation
As a technique closely aligned with real-world representation, 3D generation holds substantial promise across various sectors, including medical imaging, motion capture, asset production, and scene reconstruction, etc. However, when compared to 2D image generation, distinctive high-resolution elements such as volumetric data or point clouds present unique challenges, significantly escalating computational demands. Several efficient methodologies (Bieder et al., 2023; Zhou et al., ; Tang et al., 2023b; Park et al., 2023; Du et al., 2024; Wu et al., 2024) have been proposed, particularly concentrating on enhancing the sampling process and optimizing the architectural framework, which further handles the computational complexity inherent. One of the most prevalent approaches involves designing more efficient sampling schedules (Bieder et al., 2023; Li et al., 2024c; Yu et al., 2024b; Zhou et al., ). By utilizing larger sampling step sizes, modifying the sampling strategy between 2D and 3D, or incorporating multi-view parallelism, these techniques address the key bottlenecks in the sampling process, thereby improving sampling efficiency. Moreover, the incorporation of novel architectures, such as state-space models and lightweight feature extractors (Mo, 2024; Tang et al., 2023b), alleviates the computational burden of processing 3D data, significantly enhancing model efficiency.